# Solving Empirical Bayes via Transformers

### v1

@@ B.2 Identities on Worst Prior
where the last equality follows from that $\frac{e_{\pi_{!}}(x)}{m_{\pi_{!}}(x)}=f_{\pi_{!}}(x)$. Therefore the conclusion follows. ∎
@@ Appendix C Pseudocode on Robbins Approximation via Transformers
Appendix D Further Analysis on Experimental Results
D.1 Synthetic Experiments
We recall that our synthetic experiments are measuring regret w.r.t. sequence length for both the neural and worst-prior. In the main section, we show a plot of how the regret decreases with sequence length; here, we provide a more comprehensive result on the Plackett-Luce rankings in Table 5, along with the $p$-value by pairwise $t$-test of T18 and T24 against relevant classical baselines, as per Table 6 and Table 7. From the $p$-value we conclude that the transformers outperform other baselines by a significant margin on various experiments. (except in a handful of cases).
@@ D.1 Synthetic Experiments
Experiments GS Robbins ERM NPMLE T18 T24
Neural-128 -0.003 -3.196 0.965 4.310 7.497 7.696
Neural-256 -0.023 -3.090 1.678 4.885 7.624 8.002
Neural-512* -0.044 -3.016 2.421 5.534 7.646 8.084
Neural-1024 -0.066 -2.930 3.032 6.197 7.579 7.983
Neural-2048 -0.092 -2.813 3.430 6.806 7.455 7.816
WP-128 - -4.925 -2.434 2.476 7.416 4.945
WP-256 - -2.470 2.470 4.943 9.878 7.412
WP-512 - -2.466 2.463 4.924 9.842 7.385
WP-1024 - -2.738 2.735 8.543 8.664 5.476
WP-2048 - -2.468 2.470 9.863 7.405 4.938
Multn-512 0.505 -2.664 2.239 4.877 9.686 7.339
Multn-1024 0.463 -2.635 3.328 5.764 9.615 8.240
Multn-2048 0.471 -2.728 3.432 5.963 8.971 8.811
@@
D.2 Real Data Experiments
In the main section, we focused on reporting the RMSE. We supplement our finding on RMSE by showing a Plackett Luce ELO coefficient of RMSEs of the estimators in Table 8, which shows that the transformers are consistently ranked at the top.
Here, we also study what happens if we compare the Mean Absolute Error (MAE) of various estimators on each datapoint. Apart from the average percentage improvement of each algorithm over the MLE as per Table 9, we also display the $p$-values based on paired $t$-test of transformers vs other algorithms in Table 10, and Table 11. In Fig. 11, Fig. 12, Fig. 13 and Fig. 14, we also include the violin plots.
@@ D.2 Real Data Experiments
Dataset Robbins ERM NPMLE T18 T24
NHL -2.536 1.458 3.252 3.756 3.587
NHL (defender) -1.730 1.577 3.973 5.366 5.636
NHL (center) -2.739 0.399 2.111 3.118 3.408
NHL (winger) -3.350 0.674 2.271 3.004 2.756
MLB (batting) -2.981 2.991 5.145 6.221 8.575
MLB (pitching) -3.217 3.225 5.916 6.696 7.689
BookCorpusOpen 0.024 1.547 2.341 2.012 2.698
@@
Dataset Robbins ERM NPMLE T18 T24
Hockey (all) -20.14 $\pm$ 4.44 -0.20 $\pm$ 0.60 0.90 $\pm$ 0.58 0.76 $\pm$ 0.58 0.77 $\pm$ 0.59
Hockey (defender) -13.38 $\pm$ 4.24 1.44 $\pm$ 1.16 3.26 $\pm$ 1.07 3.66 $\pm$ 1.23 3.62 $\pm$ 1.22
Hockey (center) -41.43 $\pm$ 9.21 -0.65 $\pm$ 0.86 2.26 $\pm$ 0.82 2.87 $\pm$ 0.86 2.76 $\pm$ 0.86
Hockey (winger) -32.43 $\pm$ 7.23 -0.12 $\pm$ 0.78 1.43 $\pm$ 0.63 1.77 $\pm$ 0.67 1.68 $\pm$ 0.67
Baseball (batting) -25.97 $\pm$ 3.81 3.50 $\pm$ 0.32 5.22 $\pm$ 0.36 5.60 $\pm$ 0.34 5.53 $\pm$ 0.36
Baseball (pitching) -16.74 $\pm$ 2.19 3.45 $\pm$ 0.46 5.65 $\pm$ 0.48 5.60 $\pm$ 0.47 5.70 $\pm$ 0.44
BookCorpusOpen 28.05 $\pm$ 0.14 29.54 $\pm$ 0.10 29.65 $\pm$ 0.10 31.70 $\pm$ 0.15 27.85 $\pm$ 0.15
@@
T18 T24
Dataset MLE Robbins ERM NPMLE MLE Robbins ERM NPMLE
NHL 0.0103 6.69e-10 9.32e-05 0.868 0.0111 7.95e-10 1.33e-04 0.887
NHL (defender) 9.42e-07 6.53e-10 1.54e-05 1.78e-03 9.03e-07 6.22e-10 4.45e-05 0.013
NHL (center) 5.97e-07 3.04e-10 3.27e-08 2.48e-03 1.02e-06 3.98e-10 1.23e-07 8.20e-03
NHL (winger) 7.33e-06 7.49e-10 1.13e-05 0.0173 1.35e-05 7.98e-10 2.75e-05 0.0518
MLB batting 1.63e-24 6.01e-16 5.06e-12 6.57e-07 3.68e-24 6.74e-16 1.49e-11 3.04e-07
MLB pitching 5.67e-21 2.76e-16 6.36e-12 0.780 7.27e-22 2.38e-16 7.12e-13 0.203
BookCorpusOpen $<$1e-100 $<$1e-100 5.31e-26 3.62e-23 $<$1e-100 9.5e-69 1-6.32e-10 1-1.87e-12
Dataset Robbins ERM NPMLE T18 T24
NHL -2.928 -0.023 1.769 1.490 1.435
NHL (defender) -2.182 0.795 2.170 2.674 2.667
NHL (center) -3.201 -0.472 1.678 2.617 2.525
NHL (winger) -2.820 0.196 1.464 2.302 1.916
MLB batting -3.076 3.080 5.572 7.968 7.391
MLB pitching -3.332 3.331 6.677 6.253 7.146
BookCorpusOpen 4.325 4.896 5.067 5.319 4.704
Figure 11: Violin plots of MAE ratio achieved by multiple estimators over MLE on NHL.
Figure 12: Violin plots of MAE ratio achieved by multiple estimators over MLE on MLB batting.
Figure 13: Violin plots of MAE ratio achieved by multiple estimators over MLE on MLB pitching.
Figure 14: Violin plots of MAE ratio achieved by multiple estimators over MLE on BookCorpusOpen.

### v2

@@ B.2 Identities on worst-case prior
where the last equality follows from that $\frac{e_{\pi_{!}}(x)}{m_{\pi_{!}}(x)}=f_{\pi_{!}}(x)$. Therefore, the conclusion follows. ∎
@@ Appendix C Pseudocode on Robbins approximation via transformers
Appendix D Estimates on resources needed
We briefly detail an estimate the amount of compute we used in producing this work.
Pre-training stage. On 2 Nvidia Volta V100 GPUs and 40 CPUs, our model T24r takes about 28 hours to train, while L24r takes about 3 hours (due to its linear complexity). Together with other models trained and evaluated during hyperparameters search, we estimate the whole pretraining process to take somewhere between 1000 to 2000 GPU hours.
Synthetic experiments. Each experiment described in Section 5 is done on 2 Nvidia Volta V100 GPUs and 40 CPUs for computation (340 GB total RAM), possibly split between different tasks. The time taken ranges from a few minutes (Robbins/MLE) to a few days (NPMLE even with reduced number of batches used per prior). For non-NPMLE tasks, one of the most computationally intensive task is T24r on $n=2048$ case, which takes around 32 GPU hours. We note that it is possible to use only CPU for these tasks, although it will take longer to run.
Real experiments. Evaluation of each estimator (including NPMLE) on the sports dataset in Section 6 takes no more than a few minutes, where 2 Intel CPUs (8 GB) would suffice. Evaluation of BookCorpusOpen dataset can take a few hours (T24r) to one day (NPMLE).
Appendix E Further analysis on synthetic experiments
E.1 Performance
We recall that our synthetic experiments are measuring regret w.r.t. sequence length for both the neural and worst-prior. In the main section, we show a plot of how the regret decreases with sequence length; here, we provide a more comprehensive result on the Plackett-Luce rankings in Table 3, along with the $p$-value by pairwise $t$-test of T24r against relevant classical baselines, as per Table 4. From the $p$-value we conclude that the transformers outperform other baselines by a significant margin on various experiments. (except in a handful of cases).
@@ E.1 Performance
Experiments GS Robbins ERM NPMLE T24r L24r
Neural-128 -0.004 -3.384 0.979 4.542 7.104 7.005
Neural-256 -0.024 -3.271 1.713 5.194 7.402 7.192
Neural-512 -0.045 -3.157 2.484 5.894 7.533 7.157
Neural-1024 -0.067 -3.015 3.120 6.703 7.525 7.082
Neural-2048 -0.092 -2.834 3.461 7.515 7.487 7.009
WP-128 - -4.900 -2.429 2.475 4.942 7.407
WP-256 - -2.466 2.469 4.934 7.403 9.874
WP-512 - -2.724 2.725 5.452 8.581 8.556
WP-1024 - -2.472 2.468 4.937 9.881 7.409
WP-2048 - -2.462 2.462 4.925 9.858 7.393
Multn-128 0.462 -4.808 -2.500 3.186 5.495 7.796
Multn-256 0.443 -4.733 -2.413 3.192 5.510 7.827
Multn-512 0.502 -2.646 2.230 4.847 7.301 9.753
Multn-1024 0.455 -2.527 3.212 5.550 7.896 9.955
Multn-2048 0.471 -2.720 3.419 5.939 8.324 9.320
@@
Appendix F Further Analysis on real data experiments
F.1 Comparison with deep-learning Bayesian models
We describe why Bayesian models do not solve the real data experiments based on our EB modelling. Fundamentally, this can be seen from the following two aspects: unsupervised vs supervised, and adaptability to prior misspecification.
Supervised vs unsupervised. In EB our predictions are unsupervised at test time: we estimate $\theta_{1},\cdots,\theta_{n}$ based only on inputs $X_{1},\cdots,X_{n}$, with no prompts in the form $(X,\theta)$, and no assumption placed on $\pi$ from which $\theta$ is drawn. This is especially evident in our experimental setting in Section 6 where we make predictions $Y_{i}$’s based only on $X_{i}$’s.
On the other hand, in simulation based inference (SBI) one samples many $(\theta,X)$ pairs from the prior $\pi$. For example, [ZMSDH24] fixes a prior on $\theta$, generate many pairs $(\theta,X)$, and train a neural network to compute posterior $p_{\theta|X}$ by minimizing a suitable entropy loss. Note that in EB, such information on $\theta$ is unknown at test time, and the prior is itself unknown and has to be estimated from the input sequence $X_{1},\cdots,X_{n}$ alone.
Adaptability to prior misspecification. In EB, there is no assumption at test time on the prior from which the labels $\theta$ are drawn. In the case of transformers, for example, we evaluate them on priors unseen during the pretraining phase, (i.e. multinomial prior-on-priors and worst-case prior). Our pre-trained transformers are willing to estimate the prior at test time and adapt to it to predict $\mathbb{E}[\theta|X]$. Not only do they outperform the classical estimators: EB (NPMLE, Robbins, ERM-monotone) and non-EB (MLE, gold-standard), they also leverage longer sequences to produce better estimates (i.e. length generalization). Note that these classical EB algorithms also provably achieve decreasing regret upon longer sequences ([PW21, Theorem 2], [JPW22, Theorem 3], [JPTW23, Theorem 1] regardless of the prior (provided it fulfills certain light-tailedness condition), which is an advantage EB estimators enjoy over the non-EB estimators.
In contrast, the problem of different prior at training and test time (i.e. model misspecification) presents significant challenges in Bayesian inference, as acknowledged in [ZMSDH24, Appendix S2.8] and [CBL20, page 9 before section 3D]. The discussion of the gold-standard estimators in Appendix A.1 illustrates the fundamental limit of non-EB estimators in this respect.
Examples. Some recent works of Bayesian inference include TabPFN [HMP+25], ACE [CLH+25], Simformer [GDW+24] and Generative Bayesian Computation (GBC) [PRS24]. These are all supervised in the following sense:
  • TabPFN and ACE receive prompts containing ‘input, label’ pairs at test time, as illustrated in [HMP+25, Figure 1(a)] and [CLH+25, (1)];
  • Simformer requires simulation for each task (e.g. Appendices A2.2 and A3.2 of [GDW+24] detail the training protocols that include the number of simulations for each task)
  • GBC uses simulation (e.g. Section 2 of [PRS24] mentions the need of generative AI to simulate from all distributions).
Note that in the example of the real dataset, $\theta$ could even be impossible to obtain (for example, in the NHL or MLB dataset, $\theta$ represents the innate ability of an athlete and $X$ the number of goals scored).
We acknowledge that these ‘input-label’ pairs could have been obtained in an experimental setting different than ours. E.g. in the NHL or MLB dataset, prompts in the form of $(X_{:,j-t},X_{:,j-t+1})$ representing a player’s score in seasons $j-t$ and $j-t+1$ can be obtained, where $t=1,2,\cdots,T$ are lookback windows (thus here the label is not $\theta$ but an independent set of observations). These and $X_{:,j}$ can be given to these deep-learning Bayesian models before asking them to predict $X_{:,j+1}$. Note that in our EB setting, $T=0$, which these frameworks are incapable of. Constructing EB estimators to leverage extra information when $T\geq 1$ is beyond the scope of this paper.
F.2 More comparisons on the RMSE metric
[…]
@@ F.2 More comparisons on the RMSE metric
Dataset Robbins ERM NPMLE T24r L24r
NHL (all) -2.511 1.453 3.348 3.844 4.805
NHL (defender) -1.730 1.617 3.511 4.222 4.720
NHL (center) -2.710 0.439 2.120 3.020 3.197
NHL (winger) -3.308 0.654 2.236 3.221 3.814
MLB (batting) -3.675 0.862 3.994 6.995 5.257
MLB (pitching) -3.466 0.543 4.273 7.520 6.166
BookCorpusOpen -0.018 1.523 2.312 1.610 1.558
[…]
Dataset MLE Robbins ERM NPMLE
NHL (all) 1.51e-07 3.43e-11 1.96e-08 4.54e-04
NHL (defender) 1.05e-08 1.61e-10 3.46e-08 0.00225
NHL (center) 8.23e-08 1.79e-11 4.87e-08 2.95e-06
NHL (winger) 9.97e-08 7.36e-13 1.71e-07 3.92e-04
MLB-Bat 6.57e-17 9.89e-19 2.42e-11 1.20e-06
MLB-Pitch 1.38e-15 3.30e-20 3.66e-13 7.26e-09
BookCorpusOpen $<$ 1e-100 $<$ 1e-100 1 - 4.53e-4 $>$ 1- 1e-100
[…]
Dataset MLE Robbins ERM NPMLE
NHL (all) 1.01e-06 2.07e-11 9.38e-07 2.35e-04
NHL (defender) 6.47e-08 5.95e-11 4.84e-07 0.00203
NHL (center) 1.03e-06 9.78e-12 4.89e-07 6.05e-04
NHL (winger) 1.57e-07 6.74e-13 5.09e-07 7.76e-04
MLB batting 2.78e-16 9.79e-19 7.28e-11 1.61e-04
MLB pitching 1.66e-15 3.31e-20 1.15e-12 1.75e-06
BookCorpusOpen $<$ 1e-100 $<$ 1e-100 8.06e-95 1.76e-08
F.3 Evaluation on the MAE metric
We compare the Mean Absolute Error (MAE) of various estimators to supplement the RMSE results. Apart from the average percentage improvement of each algorithm over the MLE as per Table 9, we also display the $p$-values based on paired $t$-test of transformers vs other algorithms in Table 11, and Table 10. In Fig. 6(a), Fig. 6(b), Fig. 6(c) and Fig. 6(d), we also include the violin plots.
@@ F.3 Evaluation on the MAE metric
Dataset Robbins ERM NPMLE T24r L24r
NHL (all) -20.14 $\pm$ 4.44 -0.20 $\pm$ 0.60 0.88 $\pm$ 0.57 0.93 $\pm$ 0.53 1.12 $\pm$ 0.67
NHL (defender) -13.38 $\pm$ 4.24 1.44 $\pm$ 1.16 3.26 $\pm$ 1.07 3.09 $\pm$ 1.10 3.64 $\pm$ 1.33
NHL (center) -41.43 $\pm$ 9.21 -0.65 $\pm$ 0.86 2.26 $\pm$ 0.82 2.80 $\pm$ 0.80 2.94 $\pm$ 0.96
NHL (winger) -32.43 $\pm$ 7.23 -0.12 $\pm$ 0.78 1.43 $\pm$ 0.63 1.82 $\pm$ 0.62 2.14 $\pm$ 0.67
Baseball (batting) -49.64 $\pm$ 4.74 0.49 $\pm$ 0.24 1.57 $\pm$ 0.21 1.72 $\pm$ 0.20 1.85 $\pm$ 0.20
Baseball (pitching) -38.16 $\pm$ 2.87 0.08 $\pm$ 0.21 1.55 $\pm$ 0.19 1.76 $\pm$ 0.18 1.70 $\pm$ 0.17
BookCorpusOpen 28.05 $\pm$ 0.14 29.54 $\pm$ 0.10 29.65 $\pm$ 0.10 23.17 $\pm$ 0.21 29.72 $\pm$ 0.12
Dataset Robbins ERM NPMLE T24r L24r
NHL (all) -2.849 0.078 1.844 1.680 2.268
NHL (defender) -2.181 0.810 2.423 2.104 2.679
NHL (center) -3.139 -0.401 1.556 2.546 2.654
NHL (winger) -2.817 0.198 1.356 2.334 2.842
MLB batting -3.909 0.872 4.309 4.645 6.386
MLB pitching -3.977 0.002 3.461 6.627 5.166
BookCorpusOpen 3.585 4.191 4.395 2.964 4.131
@@
Dataset MLE Robbins ERM NPMLE
NHL (all) 0.00124 6.31e-10 2.83e-06 0.195
NHL (defender) 2.35e-06 1.36e-09 0.000253 0.929
NHL (center) 1.35e-07 3.1e-10 1.12e-08 0.000236
NHL (winger) 1.33e-06 6.63e-10 1.88e-06 4.58e-06
MLB-Bat 1.63e-16 2.39e-19 4.60e-10 0.00662
MLB-Pitch 4.73e-18 4.38e-20 1.25e-14 1.85e-06
BookCorpusOpen $<$ 1e-100 $>$ 1- 1e-100 $>$ 1- 1e-100 $>$ 1- 1e-100
[…]
Dataset MLE Robbins ERM NPMLE
NHL (all) 0.00189 4.42e-10 2.04e-06 0.0106
NHL (defender) 3.88e-06 4.13e-10 9.28e-05 0.0337
NHL (center) 1.56e-06 2.89e-10 7.49e-08 0.00198
NHL (winger) 3.55e-07 5.37e-10 3.88e-07 1.04e-06
MLB batting 3.92e-17 2.10e-19 8.51e-11 5.43e-05
MLB pitching 1.84e-18 4.61e-20 1.52e-14 6.20e-04
BookCorpusOpen $<$1e-100 1.43e-55 2.63e-26 2.18e-16
(a) NHL Hockey
(b) MLB Batting
(c) MLB Pitching
(d) BookCorpusOpen.
Figure 6: Violin plots of MAE ratio of ERM-monotone (blue), NPMLE (orange), T24r (green) and L24r (red) over MLE over various datasets. Note that T24r and L24r show a general improvement for the MLB datasets, but for bookcorpus they are generally affected by the outliers. The horizontal line at 1.0 indicates the threshold of MLE’s RMSE. Robbins is not shown due to the wide variance.